OpenStudy (anonymous):

Statistics Help: assume that a random sample is used to estimate a population proportion p. Find the margin of error E that corresponds to the given statistics and confidence level n=600, x=240, 90% confidence

OpenStudy (anonymous):

A \((1-\alpha)\times100\%\) confidence interval for a proportion has the form \[\left(\hat p-Z_{\alpha/2}\sqrt{\frac{\hat p(1-\hat p)}{n}},~\hat p+Z_{\alpha/2}\sqrt{\frac{\hat p(1-\hat p)}{n}}\right)\] where \(\hat p\) is the estimated proportion, \(Z_{\alpha/2}\) is the critical value used for the significance level \(\alpha\), and \(n\) is the sample size. The error margin is the second term, \[MOE=Z_{\alpha/2}\sqrt{\frac{\hat p(1-\hat p)}{n}}\]